A B  C D A B  CD  E  FG H E FG H A B  C E F G  H

Fermat’s Theorem Contradicted? :

9^{3} + 10^{3} = 12^{3} ??

From Dissections: Plane & Fancy, by Greg Frederickson, Cambridge University Press, to be published October 1997.

 See http://www.cs.purdue.edu/homes/gnf/book.html